OpenStudy (anonymous):

the inverse function of f(x)=5x-9 and g(x)=(x+5)/9

OpenStudy (jdoe0001):

flip the letters to get the inverse

OpenStudy (jdoe0001):

\(\bf f(x)=5x-9 \implies \color{red}{y} =5\color{blue}{x}-9 \qquad \qquad f^{-1}(x) \implies \color{blue}{x} = 5\color{red}{y}-9\)

OpenStudy (debbieg):

Huh?? no no no. You do need to write in "y=..." notation, rather than "f(x)". Then you SWITCH the x and the y (although some books may teach is slightly differently, but I'll go with this method for now because I think it is most standard). That is, you write an x in where the y is, and write in y's for any x's. THEN you SOLVE for y. So, for: \(\Large f(x)=5x-9 \) you write with a y: \(\Large y=5x-9 \) Switch the y for x and x for y: \(\Large x=5y-9 \) but THIS is NOT the inverse function. Now SOLVE for y: \(\Large x+9=5y \) \(\Large y=\dfrac{x+9}{5} \) THAT's your inverse: \(\Large f^{-1}(x)=\dfrac{x+9}{5} \) This is usually called the "switch and solve" method. Now see if you can do it with your other function. :)

OpenStudy (anonymous):

thanks DebbieG